2. Covered interest parity interest parztu We would lke to buy Polish zloty (PLN) delivered in one year from now. The PLN/USD exchange rate is 3.7738 and the NOK/USD exchange rate is 8.5765. One-year fixed borrowing and deposit rates are 2.46% and 2.04% in Norway and Poland, respectively. (a) What would a reasonable forward rate be? (b) What would your trading strategy be if a bank had offered you a forward contract for delivery of PLN in one year at rate NOK/PLN 2.2500?
Answer #1

1) As per covered interest rate parity

Formula

(1 + interest rate (base currency)) = (F/S) * (1+ interest rate (foreign currency)

So F = S * ((1+ Interest of base) / ( 1 + interest of foreign)

Calculation of Spot rate

= PLN/USD = 3.7738 and NOK/USD = 8.5765

therefore, PLN/NOK = 3.7738/8.5765 = 0.4400

Forward rate of PLN in terms of NOK = 0.4400 * (1+0.0246)/(1+0.0204) = 0.4418

(Just for informaton : Spot rate of NOK in terms of PLN = 8.5765/3.7738 = 2.2726

Forward rate for NOK in terms of PLN = 2.2726*(1+0.0204)/(1+0.0246) = 2.2633 )

2) If bank has offered a forward contract delivery of PLN in one year at the rate of NOK/PLN 2.2500 then we should

1. Borrow the PLN now from bank and then convert it in NOK

2. Deposit the NOK to take interest which is higher than in Poland

3. After one year convert the currency in PLN into NOK.

As there is a difference between bank rate and the forward rate that should be prevailing in the market as per covered interest rate parity, There would be arbitrage interest prevailing in the market.

In short we can borrow from bank POL and we have to pay it after one year. These POL will be converted into NOK to earn higher interest and when the market rate for POL would be higher than the bank rate we will pay the POL price to the bank.